OpenStudy (anonymous):

Find the unit rate. The tub filled with 12 gallons of water in 5 minutes.

OpenStudy (math&ing001):

12 gallons in 5 min ? gallons in 1 min the rule of 3: 12*1/5 = 2.4gallons per minute

OpenStudy (anonymous):

oh my gosh!! thanks @math&ing001

OpenStudy (math&ing001):

Your much welcome :)

OpenStudy (texaschic101):

I usually set these up as proportions... 12/5 = x/1 (12 gal to 5 min = x gal to 1 min) cross multiply 5x = 12 x = 12/5 x = 2 2/5 gallons or 2.4 gallons either way you do it, you will get the same answer :)

OpenStudy (anonymous):

If it takes 5 minutes to fill the tub of water with 12 gallons, the answer is 2.4 gallons per minute. I came to this conclusion because 1/5 (for the minutes)*12=2.4.